A Bibliography of Publications on BMDP

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

17 November 2014 Version 1.10

Title word cross-reference

1985 [All86].

4V [DT79].

accompany [BMD89, BMD90, BMD92, Dix88, Dix90, Dix92]. ACO [SG80]. additions [Dix83]. advanced [Chu83, Chu84, Chu85]. algorithms [Chu83, Chu84, Chu85]. analisis [SM82]. Analyse [Laf79, Laf81, Ber77]. analyses [SC87]. Analysis [DEL83, KW85, Bro86, DT79, Mic70]. Annotated [SG80]. Annotations [Bar83]. applications [Chu83, Chu84, Chu85, Chu86, Chu87, Laf81]. Applied [SC87]. approaches [Fra78a]. Auswertungen [KW85].

BALANOVA [DEL83]. base [Chu86, Chu87]. Berkeley [Uni83]. between [Fra79b]. biologiques [Laf79, Laf81]. Biomedical [DB79, Uni77, Dix64, Dix65, Dix67, Dix70b, Dix70a, Dix73, Dix75, DB77]. BMD [Uni77, Chu83, Chu84, Chu85, Chu86, Chu87, Dix64, Dix65, Dix67, Dix70b, Dix70a, Dix73, Fra76, Fra78a, Laf79, Laf81, Mic70]. BMDP [Uni83, Bar83, Ber77, BD83, Bro86, DT79, DEL83, Dix75, DB77, DB79, Dix81, Dix83, DB85, Dix88, Dix90, Dix92, DCT89, EN88, Fra76, Fra78a, Fra78b, Fra79a, Fra79b, Hil79, Hil82, Hil84, KW85, Laf79, Laf81, SM82, SG80, SC87, TF80]. BMDP-77 [DB77, Fra79b]. BMDP-79 [DB79]. BMDP-V [SG80].

Canonical [Mic70]. Categorical [Bro86]. CLISTs [TF80]. CMS [Uni83].

Comparaison [Ber77]. Comparative [DEL83]. comparison [EN88]. components [SG80]. Computer [All86, Bar83, Fra76, Uni77, Dix64, Dix65, Dix67, Dix70b, Dix70a, Dix73, Dix75, DB77, DB79, Hil82, Hil84, SG80, All86]. condensed [Hil82, Hil84]. contingency [Mic70]. correlation [Mic70].

data [Bro86, Chu86, Chu87, Fra78a, Fra78b, KW85]. Database [Chu83, Chu84, Chu85]. Datenanalysessyteme [KW85]. datos [SM82]. del [SM82]. Design [Bar83]. developments [Fra79b]. digest [BMD89, BMD90, BMD92, Hil82, Hil84]. Distributed [Chu86, Chu87]. d’ordinateur [Laf79].

Einführung [KW85]. elementary [DCT89]. estadisticos [SM82]. Examples [DEL83].

features [Fra79b]. February [Chu83, Chu84]. file [Fra78b]. final [Chu83, Chu84]. Fortran [Fra79a]. future [Bro86].

general [DT79]. GENSTAT [Ber77]. geograficos [SM82]. get [Uni83]. guide [Fra79a, Hil82, Hil84].

handbook [SC87]. Hotel [All86].

IBM [TF80]. II [Fra79b]. Input [Bar83]. interactively [TF80]. interest [Fra79b]. Interface [All86, Fra79b]. Introduccion [SM82]. Introduction [KW85].

January [Chu83, Chu84].

Kentucky [All86].

Lexington [All86]. logiciels [Ber77, Laf81].

management [Chu83, Chu84, Chu85, Chu86, Chu87]. manual [Dix88, Dix90, Dix92]. March [All86]. matricielle [Laf79]. matrix [Fra78b]. MINITAB [DCT89]. Modern [DCT89]. multivariate [DT79]. multivariée [Laf79, Laf81]. MVS [TF80]. MVS/TSO [TF80].

Novel [Mic70].

Output [Bar83, SG80]. overview [Fra79a].

P [DB77, DB79]. P-series [DB77, DB79]. Packages [Bar83]. paquete [SM82]. Part [Fra79b]. PC [EN88]. period [Chu83, Chu84]. Plaza [All86]. pour [Ber77, Laf81]. present [Bro86]. présentation [Laf79]. printing [Dix83]. probability [DCT89]. Proceedings [All86]. program [Fra78b]. programas [SM82]. programmer [Fra79a]. programmes [Laf79]. programming [DCT89]. Programs [Fra76, Uni77, BMD89, BMD90, BMD92, Dix64, Dix65, Dix67, Dix70b, Dix70a, Dix73, Dix75, DB77, DB79, Hil82, Hil84, Mic70].

quick [BMD89, BMD90, BMD92].

Radisson [All86]. real [Chu86, Chu87]. real-time [Chu86, Chu87]. reference [BMD89, BMD90, BMD92]. release [BMD89, BMD90, BMD92, Dix88, Dix90, Dix92]. report [Chu83, Chu84]. Research [Bar83]. running [TF80].

Sample [Fra78b]. sampler [Hil79]. SAS [Bar83, DCT89, EN88, Fra79b, KW85, Laf81]. science [All86]. Sciences [All86, Laf79, Laf81]. Series [Fra76, Dix70a, DB77, DB79]. Seventeenth [All86]. sigma [SG80]. sociales [Laf79, Laf81]. software [BD83, Dix81, Dix83, DB85, Dix88, Dix90, Dix92]. some [Fra79b]. special [Fra79b]. specialisés [Ber77]. SPSS [Bar83, Ber77, DEL83, EN88, KW85, Laf79, Laf81]. SPSSX [Bar83]. started [Uni83]. Statistical [Bar83, Fra76, KW85, BD83, Dix81, Dix83, DB85, Dix88, Dix90, Dix92, DCT89]. statistics [All86, DCT89, SC87]. statistique [Ber77]. statistische [KW85]. supplement [Dix70a]. Symposium [All86]. Systems [KW85].

tables [Mic70]. time [Chu86, Chu87]. trois [Ber77]. TSO [TF80].

unbalanced [Fra78a]. univariate [DT79]. usage [Laf79]. user [BMD89, BMD90, BMD92, Hil82, Hil84]. users [Fra79b]. uses [Mic70]. utilizacion [SM82].

V [SG80]. Variance [DEL83, DT79, SG80].

write [Fra78b].

X [Dix70a]. X-series [Dix70a].

References

Allen:1986:CSS

[All86] David M. Allen, editor. Computer science and statistics: Proceedings of the Seventeenth Symposium on the Interface of Computer Sciences and Statistics, Radisson Plaza Hotel, Lexington, Kentucky, March 17–19, 1985. North-Holland, Amsterdam, The Netherlands, 1986. ISBN 0-444-70018-8 (paperback). LCCN QA276.4 .S95 1985. URL http:

//catalog.hathitrust.org/api/
volumes/oclc/13455449.html;
http://www.dtic.mil/dtic/tr/

fulltext/u2/a166019.pdf;   http:
//zbmath.org/?q=an:0647.62008.

Barcikowski:1983:CPR

[Bar83] Robert S. Barcikowski. Computer Packages and Research Design: With Annotations of Input and Output from the BMDP, SAS, SPSS, and SPSSX Statistical Packages. University Press of America, Lanham, MD, USA, 1983. ISBN 0-8191-3494-5 (v. 1: paperback), 0-8191-3495-3 (v. 2: paperback), 0-8191-3496-1 (v. 3: paperback). xiv + 781 pp. LCCN QA276.4.C576 1983.

Brown:1983:BSS

[BD83] M. B. Brown and Wilfrid Joseph Dixon. BMDP statistical software. Berkeley, CA, USA, 1983 print. with additions edition, 1983. x + 733 pp. Includes index.

Bernard:1977:CTL

[Ber77] G. Bernard. Comparaison de trois logiciels specialisés pour l’analyse statistique: GENSTAT, BMDP, SPSS. Technical report, Departement de mathématique, Université d’Orsay, Orsay, France, 1977. 15 pp.

BMDP:1989:BUD

[BMD89] BMDP Statistical Software (Firm). BMDP user’s digest: quick reference for the BMDP programs: to accompany the 1988 software release. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, 1989. ISBN 0-935386-14-9. 218 pp. LCCN W 26.5 B661 1989.

BMDP:1990:BUD

[BMD90] BMDP Statistical Software (Firm).

BMDP user’s digest: quick reference for the BMDP programs: to accompany the 1990 software release. Berkeley, CA, USA, 1990. 233 pp.

BMDP:1992:BUD

[BMD92] BMDP Statistical Software (Firm).

BMDP user’s digest: quick reference for the BMDP programs: to accompany BMDP release 7. Berkeley, CA, USA, 1992. 251 pp. Developed by BMDP Statistical Software, Inc. Includes index.

Brown:1986:CDA

[Bro86] Morton B. Brown. Categorical data analysis in BMDP: present and future. In Allen [All86], pages 257–262. ISBN 0-444-70018-8 (paperback). LCCN QA276.4 .S95 1985. US$65. URL http:

//catalog.hathitrust.org/api/
volumes/oclc/13455449.html;
http://www.dtic.mil/dtic/tr/
fulltext/u2/a166019.pdf;   http:
//zbmath.org/?q=an:0647.62008.

Chu:1983:DMA

[Chu83] Wesley W. Chu. Database management algorithms for advanced BMD applications: final report for the period from February 1, 1982 to January 31, 1983. Los Angeles, CA, USA, 1983. (various) pp. “Prepared for Ballistic Missile Defense Advanced Technology Center.” Includes bibliographies.

Chu:1984:DMA

[Chu84] Wesley W. Chu. Database management algorithms for advanced BMD applications: final report for the period from February 1, 1983 to January 31, 1984. Los Angeles, CA, USA, 1984. (various) pp. “Prepared for Ballistic Missile Defense Advanced Technology Center.” Includes bibliographies.

Chu:1985:DMA

[Chu85] Wesley W. Chu. Database management algorithms for advanced BMD applications. Los Angeles, CA, USA, 1985. 103 pp.

Chu:1986:DDB

[Chu86] Wesley W. Chu. Distributed data base management for real-time BMD applications. Los Angeles, CA, USA, 1986. (various) pp.

Chu:1987:DDB

[Chu87] Wesley W. Chu. Distributed data base management for real-time BMD applications. Los Angeles, CA, USA, 1987. (various) pp. Includes bibliographical references.

Dixon:1977:BBC

[DB77] Wilfrid Joseph Dixon and M. B. Brown. BMDP-77, biomedical computer programs, P-series. Berkeley, CA, USA, 1977. xiii + 880 pp. First published in 1975 under title: BMDP: biomedical computer programs.

Dixon:1979:BBC

[DB79] Wilfrid Joseph Dixon and M. B. Brown. BMDP-79: Biomedical computer programs P-series. University of California Press, Berkeley, CA, USA, 2d print., rev. and updated edition, 1979. ISBN 0-520-03569-0. xiii + 880 pp. LCCN RA409.5 .C33 1979, Z699.5 M39 U6 1979. Includes index.

Dixon:1985:BSS

[DB85] Wilfrid Joseph Dixon and M. B.

Brown. BMDP statistical software. Berkeley, CA, USA, 1985 print edition, 1985. x + 733 pp. Includes index. Cover title: BMDP statistical software manual.

Dudewics:1989:MEP

[DCT89] Edward J. Dudewics, Pinyuen Chen, and Baldeo K. Taneja. Modern elementary probability and statistics, with statistical programming in SAS, MINITAB, BMDP. American Sciences Press, Syracuse, NY, USA, 1989. ISBN 0-935950-19-2. ix + 375 pp. LCCN QA273 .D793 1989. US$34.50. Includes index. Includes bibliographical references.

DeBoer:1983:AVC

[DEL83] Donald P. DeBoer, Frederick T.

Evers, and Douglas Garfield Link. Analysis of Variance: Comparative Examples, BALANOVA, BMDP and SPSS. Social Science Computing Laboratory, University of Western Ontario, London, Ontario, Canada, 1983. ISBN 0-7714-0427-1. (various) pp.

Dixon:1964:BBC

[Dix64] Wilfred Joseph Dixon. BMD:

biomedical computer programs. Health Sciences Computing Facility, Department of Preventive Medicine and Public Health, School of Medicine, University of California, Los Angeles, 1964. vi + 585 pp. LCCN RA409.5 .C33 1964. Includes bibliographies.

Dixon:1965:BBC

[Dix65] Wilfrid Joseph Dixon. BMD: biomedical computer programs. Health Sciences Computing Facility, Department of Preventive Medicine and Public Health, School of Medicine, University of California, Los Angeles, rev edition, 1965. viii + 620 pp. LCCN RA409.5 .C33 1965. Work under grant FR-3 of the Division of Research Resource Facilities of the National Institutes of Health. Includes bibliographies.

Dixon:1967:BBC

[Dix67] Wilfred Joseph Dixon. BMD; biomedical computer programs. Berkeley, CA, USA, [2d ed.] edition, 1967. x + 600 pp. Includes bibliographies.

Dixon:1970:BBCb

[Dix70a] Wilfred Joseph Dixon. BMD; biomedical computer programs. Xseries supplement. Berkeley, CA, USA, 1970. 260 pp.

Dixon:1970:BBCa

[Dix70b] Wilfrid Joseph Dixon. BMD; biomedical computer programs. University of California publications in automatic computation no. 2. University of California Press, Berkeley, CA, USA, [2d ed., rev.] edition, 1970. x + 600 pp. LCCN Z699.5 M39 U58 1970. Includes bibliographies.

Dixon:1973:BBC

[Dix73] Wilfrid Joseph Dixon. BMD; biomedical computer programs. University of California Press, Berkeley, CA, USA, third edition, 1973. ISBN 0-520-02426-5. vii + 773 pp. LCCN RA409.5 .C33 1973, Z699.5 M39 B13 1973. “This publication replaces numbers 2 and 3 of the series, Publications in automatic computation.” Includes bibliographical references.

Dixon:1975:BBC

[Dix75] Wilfrid Joseph Dixon. BMDP:

biomedical computer programs. University of California Press, Berkeley, CA, USA, 1975. xi + 792 pp. LCCN RA409.5.C33 1975. US$10.00. Edition of 1977 published under title: BMDP-77, biomedical computer programs, P-series.

Dixon:1981:BSS

[Dix81] Wilfrid Joseph Dixon. BMDP statistical software: 1981. Berkeley, CA, USA, 1981 ed edition, 1981. x + 725 pp. Includes index. On cover: Dept. of Biomathematics, University of California, Los Angeles.

Dixon:1983:BSS

[Dix83] Wilfrid Joseph Dixon. BMDP statistical software: 1983 printing with additions. University of California Press, Berkeley, CA, USA, 1983. ISBN 0-520-04408-8. x + 733 pp. LCCN RA409.5 .B57 1983. Includes index.

Dixon:1988:BSS

[Dix88] Wilfrid Joseph Dixon. BMDP statistical software manual: to accompany the 1988 software release. Berkeley, CA, USA, 1988. Includes index.

Dixon:1990:BSS

[Dix90] Wilfrid Joseph Dixon. BMDP statistical software manual: to accompany the 1990 software release. Berkeley, CA, USA, 1990. xviii + 1385 pp. Two volumes. Includes index. Includes bibliographical references.

Dixon:1992:BSS

[Dix92] Wilfrid Joseph Dixon. BMDP statistical software manual: to accompany BMDP release 7. California/Princeton Fulfillment Services [distributor], Princeton, NJ, USA, 1992. ISBN 0-520-08141-2 (set), 0-520-08138-2 (v. 1), 0-520-08139-0 (v. 2), 0-520-08140-4 (v. 3). (various) pp. LCCN QA276.4 .B58 1992. US$72.50 (set), US$33.00 (v. 1), US$33.00 (v. 2), US$22.00 (v. 3). Volume 3 has a distinctive title: The data manager. Includes bibliographical references and indexes.

Davidson:1979:BUM

[DT79] Michael L. Davidson and Jerome D. Toporek. BMDP 4V–general univariate and multivariate analysis of variance. Technical report / BMDP Statistical Software; no. 67. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, rev edition, 1979. 117 pp. LCCN CM.2.2.05. Includes bibliography.

Evans:1988:CSP

[EN88] Gerald Evans and William A. Newman. A comparison of SPSS PC, SAS PC, and BMDP PC. Collegiate Microcomputer, 6(2):97–106, May 1988. CODEN CMICDL. ISSN 0731-4213.

Frane:1976:BBS

[Fra76] J. W. Frane. The BMD and BMDP series of statistical computer programs. Communications of the ACM, 19:570–576, October 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frane:1978:BBA

[Fra78a] James W. Frane. BMD and BMDP approaches to unbalanced data. Technical report / Health Sciences Computing Facility, University of California, Los Angeles; no. 41. Health Sciences Computing Facility, Department of Preventive Medicine and Public Health, School of Medicine, University of California, Los Angeles, 1978. 10 pp. LCCN CM.2.2.02. Includes bibliography.

Frane:1978:SPW

[Fra78b] James W. Frane. Sample program to write a data matrix on a BMDP file. Technical report / Health Sciences Computing Facility, University of California, Los Angeles; no. 42. Health Sciences Computing Facility, Department of Preventive Medicine and Public Health, School of Medicine, University of California, Los Angeles, 1978. 5 pp. LCCN CM.2.2.02. Includes bibliography.

Frane:1979:BFP

[Fra79a] James W. Frane. The BMDP Fortran programmer’s guide: an overview. Technical report / BMDP Statistical Software; no. 66. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, 1979. 6 pp. LCCN CM.2.2.05.

Frane:1979:PII

[Fra79b] James W. Frane. Part I–The 1978 interface between BMDP-77 and SAS and some BMDP features of interest to SAS users. Part II– Recent developments in BMDP of special interest to SAS users. Technical report / BMDP Statistical Software; no. 40. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, 1979. 4 + 2 + 8 pp. LCCN CM.2.2.02. Includes bibliography.

Hill:1979:BS

[Hil79] Mary Ann Hill. BMDP sampler. Technical report / BMDP Statistical Software; no. 53. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, 1979. 93 pp. LCCN CM.2.2.03. Includes “Complete list of BMDP programs.”.

Hill:1982:BUD

[Hil82] Mary Ann Hill. BMDP user’s digest: a condensed guide to the BMDP computer programs. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, 2nd ed edition, 1982. ISBN 0-935386-02-5. 157 pp. LCCN W 26.5 H647b 1982.

Hill:1984:BUD

[Hil84] Mary Ann Hill. BMDP user’s digest: a condensed guide to the BMDP computer programs. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, 3rd ed edition, 1984. 177 pp. LCCN QA276.4 .H54 1984, W 26.5 H647b 1984.

Kuffner:1985:DSA

[KW85] Helmuth Kuffner and Reinhard Wittenberg. Datenanalysessyteme fur statistische Auswertungen: e. Einführung in SPSS, BMDP und SAS [English: Data Analysis Systems for Statistical ??: An Introduction to SPSS, BMDP, and SAS]. G. Fischer, Stuttgart, Germany, 1985. 292 pp.

Laforge:1979:AMS

[Laf79] Hubert Laforge. Analyse multivariée: en sciences sociales et biologiques: présentation matricielle et usage des programmes d’ordinateur BMD, BMDP et SPSS. Presses de l’Université Laval, Québec, Québec, Canada, 1979. vi + 450 pp.

Laforge:1981:AMP

[Laf81] Hubert Laforge. Analyse multivariée: pour les sciences sociales et biologiques avec applications des logiciels BMD, BMDP, SPSS, SAS. Etudes vivantes, Saint-Lambert, Québec, 1981. ISBN 2-7607-0075-5. xiii + 317 pp.

Mickey:1970:NUB

[Mic70] M. R. Mickey. Novel uses of BMD programs. I. Canonical correlation analysis of contingency tables. Technical report / Health Sciences Computing Facility, University of California, Los Angeles; no. 2. Health Sciences Computing Facility, Department of Preventive Medicine and Public Health, School of Medicine, University of California, Los Angeles, 1970. 8 pp. LCCN CM.2.2.01. Includes bibliography.

Snell:1987:ASH

[SC87] E. J. Snell and D. R. Cox. Applied statistics: a handbook of BMDP analyses. Chapman and Hall, Ltd., London, UK, 1987. ix + 171 pp. LCCN QA276 .C66531 1981 Suppl. “A complement to Applied statistics: principles and examples [by] D. R. Cox and E. J. Snell.” Includes index. On t.p. the registered trademark symbol “TM” is superscript following “BMDP” in the title.

Searle:1980:ACO

[SG80] Shayle R. Searle and B. A. Grimes. Annotated computer output for variance components (ACO [sigma]2): BMDP-V. Paper … in the Biometrics Unit; no. BU-716-M. Biometrics Unit, Cornell University, 1980. xxiv + 45 pp.

Santos-Preciado:1982:IUP

[SM82] Jose Miguel Santos Preciado and Carmen Muguruza Canas. Introduccion, a la utilizacion del paquete de programas estadisticos BMDP en el analisis de datos geograficos. Departamento de Geografia Humana, Facultad de Geografia e Historia, Universidad Complutense, Madrid, Spain, 1982. ISBN 84-7491-051-X. 189 pp. LCCN G70.2 .S26 1982.

Toporek:1980:IMC

[TF80] Jerome D. Toporek and James W.

Frane. IBM MVS/TSO CLISTs for running BMDP interactively. Technical report / BMDP Statistical Software; no. 78. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, 1980. 11 pp.

BMDP:1977:BBC

[Uni77] University of California, Los Angeles. Health Sciences Computing Facility. BMD, biomedical computer programs. University of California Press, Berkeley, CA, USA, 1977. 67–764 pp. Includes bibliographies.

BMDP:1983:BUC

[Uni83] University of California, Berkeley.

Computing Services. BMDP under CMS: how to get started with BMDP at Berkeley. Computing Services, University of California, 1983. 12 pp. LCCN CM.3.2.01. Includes bibliography.